Find an antiderivative for each function. Check your answers by differentiation. = (iv) 1-8csc 2x (v) 3 sec 3x tan 3x (0)
Answer #1

Antiderivative of following functions 4 312 3/2 | Irrida - ne la Bežti ztite=2 C=u +c=2 ntc 3/2 using result fondre a lot ...lig Judech s Joten 16 = 25+6. Antidestinative of the r = 255 +6. Verification So derivative of a (2544(2) = 20 (JE) + a 1 (2)